OpenStudy (anonymous):

solve the given equation 7^4n-7

myininaya (myininaya):

this is not an equation

OpenStudy (anonymous):

but it is.....everything after the 7 is a power

myininaya (myininaya):

an equation has an =

myininaya (myininaya):

so what do you want to do

OpenStudy (anonymous):

it wants to knkow what n equals

myininaya (myininaya):

is it \[7^{4n}-7=0\]

OpenStudy (anonymous):

no its 4n-7

myininaya (myininaya):

\[7^{4n-7}=0\]

OpenStudy (anonymous):

the options i have are 11/4, -3, -3/4, or -1

myininaya (myininaya):

exponential functions are never zero so what goes after the equal sign?

OpenStudy (anonymous):

2401

myininaya (myininaya):

ok thank you that part is important lol

OpenStudy (anonymous):

sorry i didnt even see that

myininaya (myininaya):

\[7^{4n-7}=2401\] now we have an equation

myininaya (myininaya):

\[7^{4n-7}=7^4\]

myininaya (myininaya):

we have the same base so we can set 4n-7=4 add 7 to both sides 4n=4+7 4n=11 n=11/4

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

3^n-12=1/81